T-474 phone conference

Thursday 29 june 2006, 17:00 GMT+1 (09:00 PDT)

Phone number inside US : 866-800-0606
Phone number from outside US : (+ 1) 973-532-4913
Meeting code : 98195

general points

  • Email list set up, info see mail
  • Plots and analysis progress can be found as usual in the T474 analysis eLog
  • Reprocessing of external clock basically done, files can be found at this location, the runlist with more details is still in the eLog, Finished runs so far :
    367 - 501, 526 - 536, 679 - 705 ( all except 960 - )
    with all BPMs calibrated. Note for runs 960- no BPM 31,32 calibration will exist

discussion points

  • Frequencies, gammas and other quantities vs time (Chit and Oleg)
  • Phase dependence of the amplitude in the DDC ( see elog entries from mark and bino )
    • relative misalignments between BPMs in each triplet
    • relative misaligmnents between triplets
  • General status roundup before next run + runplan for July run (some ideas) :
    • first shift to check freqs, gammas etc
    • aligment checks with new data !
    • a shift just calib - resolution run 20k - calib - resolution run ....
    • long "corrector scan", to have numbers of resolution at different offsets ?


Minutes of the 29 June 2006 T474 Phone conference

 - We need alignment numbers asap, Yury will look to some later runs
   and get those numbers today, Bino will do the same independently
   and send the numbers around tomorrow. (UK time that is :-)

 - As to the quick study that was done concerning the phase dependence of
   the amplitude :

   o results that the amplitude of the residual sin wave decreases as
     a function of the filter bandwidth is confusing, one would expect
     opposite behaviour.

     -> need to look at the amplitude of the signal relative to the
        amplitude of the residual sin wave versus filter bandwidth,
        see whether that increases, the combined info will give a
        clearer view on what's going on.

     -> Yury suggested to look at the Fourier transform of the
        downmixed signal before the filtering, so to really see the 0
        and 2-omega component.

     -> Substitute maybe the gaussian filter with a square box filter
        (which probably has

 - Fil is copying the external clock runs that were reprocessed to
   SLAC, the files will be in :


 - Fil is looking at the reprocessing internal clock runs, he got
   updated frequencies (see eLog) and is looking to minimize the

 - Oleg and Chit are having a look at the time stability of
   frequencies and gammas. Updated plots are and will be put to the
   Wiki at Berkeley.

 - Shift plan : we have probably 1/3 of the time available, so 10
   shifts. First couple of shifts should be to check out the
   frequencies, gammas and the calibrations. Then we need to take a
   lot of stability data over more than one hour, say just 2 shifts in
   a row or so, stability over 16 hours or so... We will reiterate
   this via mail and than send it to Mike for his runplan...
back to main